OpenStudy (anonymous):

Find h(g(2)) g(x)=-3x^2+7 and h(x)=square root x-2 how do i find this, confused?

OpenStudy (jdoe0001):

\(\bf g(x)=-3x^2+7\qquad h(x)=\sqrt{x-2} \\ \quad \\ h({\color{red}{ x}})=\sqrt{{\color{red}{ x}}-2}\implies h(\quad {\color{red}{ g(x)}}\quad )=\sqrt{{\color{red}{ g(x)}}-2} \\ \quad \\ \implies h(\quad {\color{red}{ g(x)}}\quad )=\sqrt{({\color{red}{ -3x^2+7}})-2}\)